One rectangle has dimensions that are double those of a second rectangle. Let A be the area of the larger rectangle. Which formula would necessarily give you the area of the smaller rectangle, B?
A. b= a/8
B. b= a/4
C. b= 2a
D. b= 4a

Answer by TimothyLamb(4379)

---
small rectangle dimensions:
x = width
y = length
---
small rectangle area:
b = xy
---
large rectangle area:
a = 2x2y
a = 4xy
---
area ratio:
b/a = xy/4xy
b/a = 1/4
---
answer:
small rectangle area:
b/a = 1/4
b = (1/4)a
b = a/4
---
Solve and graph linear equations:
https://sooeet.com/math/linear-equation-solver.php
---
Solve quadratic equations, quadratic formula:
https://sooeet.com/math/quadratic-formula-solver.php
---
Solve systems of linear equations up to 4-equations 4-variables:
https://sooeet.com/math/system-of-linear-equations-solver.php